# CTPD: Cross-Modal Temporal Pattern Discovery for Enhanced Multimodal Electronic Health Records Analysis

### v2

@@ 4.1 Experimental Setup
Evaluation Metrics. The 48-hour In-Hospital Mortality (48-IHM) prediction is a binary classification with a marked label imbalance, indicated by a death-to-discharge ratio of approximately 1:6. Following previous work (Harutyunyan et al., 2019; Zhang et al., 2023b), we use AUROC, AUPR, and F1 score, for a comprehensive evaluation. The -hour Phenotype Classification (24-PHE) involves predicting the presence of 25 different medical conditions during an ICU stay, making it a multi-label classification task. For this task, we employ the AUROC, AUPR, and F1 score (Macro) for a thorough assessment of model efficacy. The F1 score threshold is determined by selecting the value that maximizes the F1 score on the validation set.
Implementation Details. We train the model with batch size of 128, learning rate of 4e-5, and Adam (Kingma and Ba, 2014) optimizer. We use a cosine annealing learning rate scheduler with a 0.2 warm-up proportion. To prevent overfitting, we implement early stopping when there is no increase in the AUROC on the validation set for 48-IHM or 24-PHENO over consecutive epochs. All experiments are conducted on 1 RTX-3090 GPU card using about 1 hour per run. We clip the norm of gradient values with 0.5 for stable training. By default, we use Bert-tiny Turc et al. (2019); Bhargava et al. (2021) as our text encoder.
Compared Methods. To ensure a comprehensive comparison, we compare our CTPD with three types baselines: MITS-only approaches, note-only approaches and multimodal approaches. For MITS-only setting, we compare CTPD with 4 baselines for imputed regular time series: RNN (Elman, 1990), LSTM (Hochreiter, 1997), CNN (LeCun et al., 1998) and Transformer Vaswani et al. (2017), and 5 baselines for irregular time series, including IP-Net (Shukla and Marlin, 2019), GRU-D (Che et al., 2018), DGM-O (Wu et al., 2021), mTAND (Shukla and Marlin, 2021), SeFT (Horn et al., 2020), and UTDE (Zhang et al., 2023b). The imputation approach follows the MIMIC-III benchmark (Harutyunyan et al., 2019). For the note-only setting, we compare our model with 6 baselines: Flat Deznabi et al. (2021), HierTrans Pappagari et al. (2019), T-LSTM Baytas et al. (2017), FT-LSTM Zhang et al. (2020a), GRU-D (Che et al., 2018), and mTAND (Shukla and Marlin, 2021). In the multimodal setting, we compare our model with 4 baselines: MMTM Joze et al. (2020), DAFT Pölsterl et al. (2021), MedFuse Hayat et al. (2022), and DrFuse Yao et al. (2024). To ensure a fair comparison, we implement Bert-tiny Bhargava et al. (2021); Turc et al. (2019) as the text encoder across all baselines. Details of baselines can be found in the Appendix B.
@@ 4.2 Comparison with SOTA Baselines
Discussion on Missing Modalities and Noisy Data Scenarios. Currently, CTPD framework is built on paired time-series and clinical notes. In practice, the dataset might have missing modalities, such as partial clinical notes or time-series data are missed. We acknowledge this is an interesting research direction. However, the study of missing modalities falls outside the scope of our work and will be explored in future research. Additionally, our evaluation is conducted on MIMIC-III Johnson et al. (2016), a large-scale, de-identified real-world dataset containing patient records from critical care units at Beth Israel Deaconess Medical Center between 2001 and 2012. Our approach is designed to be adaptable and can be seamlessly applied to other real-world databases.
@@ 6 Limitations
A limitation of our approach is its primary focus on extracting temporal semantics in the embedding space, which affects model interpretability. As a potential solution, we plan to explore retrieval-based methods Patel et al. (2024) or pretrained generative models Zhao et al. (2023) to enhance the interpretability of learned temporal pattern embeddings in future work. Besides, our work primarily focuses on exploring cross-modal temporal patterns to enhance representation learning, the scenario of missing modalities falls outside the current scope—since shared patterns across modalities cannot be fully captured in their absence. We plan to thoroughly investigate this direction in future work. Additionally, our framework is currently designed for specific clinical prediction tasks. In practice, there are various prediction tasks related to multimodal EHR analysis. Extending the proposed method into a more generalized or foundational model capable of handling multiple downstream tasks with minimal training annotations could be more practical and effective. Our future work will focus on resolving those aspects.
@@ Appendix A Experimental Setup Details
We presented the class distributions for 48-IHM and 24-PHE in Table 7 and Table 8, respectively. 48-IHM is a binary classification task, whereas 24-PHE is a multi-label classification problem with 25 labels. Notably, our 24-PHE task differs from the phenotype classification problem in the MIMIC-III benchmark but follows the setup in Zhang et al. (2023b). This setup focuses on acute care conditions that arise during ICU stays, where early prediction is critical for timely intervention. To enhance clinical relevance, we use only the first 24 hours of data for phenotype classification rather than the entire admission record. As highlighted in Yang et al. (2021), early-stage diagnosis holds greater clinical significance.
The selection of 25 phenotype labels for the 24-PHE task follows established practices in the MIMIC-III benchmark Harutyunyan et al. (2019), as also utilized in prior studies like UTDE Zhang et al. (2023b) and MedFuse Hayat et al. (2022). These labels cover conditions commonly observed in adult ICUs, including 12 critical and life-threatening conditions (e.g., respiratory failure, sepsis), 8 chronic conditions that are often considered comorbidities or risk factors (e.g., diabetes, metabolic disorders), and 5 ‘mixed’ conditions that exhibit characteristics of both chronic and acute conditions. Phenotype labels were determined using the MIMIC-III ICD-9 diagnosis table. To facilitate the translation and conversion of the abovementioned conditions, we use the Health Cost and Utilization (HCUP) Clinical Classification Software (CCS)222https://hcup-us.ahrq.gov/toolssoftware/ccs/ccs.jsp. We first mapped each ICD-9 code to its corresponding HCUP CCS category, retaining only 25 categories. Diagnoses were then linked to ICU stays using the hospital admission identifier, as ICD-9 codes in MIMIC-III are associated with hospital visits rather than specific ICU stays. To reduce label ambiguity, we excluded hospital admissions involving multiple ICU stays, ensuring each diagnosis could be associated with a single ICU stay. Please find the class distribution of 25 phenotypes in Table 8.
@@ B.1 Baselines only using MITS
CNN (LeCun et al., 1998): CNN (Convolutional Neural Network) uses backpropagation to synthesize a complex decision surface that facilitates learning.
RNN (Elman, 1990): RNN (Residual Neural Network) is trained to process data sequentially so as to model the time dimension of data.
LSTM Hochreiter (1997): LSTM (Long short-term memory) is a variant of recurrent neural network. It excels at dealing with the vanishing gradient problem and is relatively insensitive to time gap length.
Transformer (Vaswani et al., 2017): Transformer is a powerful deep learning architecture based on attention mechanism. It has great generalisability and has been adopted as foundation model in multiple research domains.
IP-Net (Shukla and Marlin, 2019): IP-Net (Interpolation-Prediction Network) is a deep learning architecture for supervised learning focusing on processing sparse multivariate time series data that are sampled irregularly.
GRU-D (Che et al., 2018): GRU-D is based on GRU (Gated Recurrent Unit). It incorporates features of missing data in EHR into the model architecture to improve prediction results.
DGM-O (Wu et al., 2021): DGM-O (Dynamic Gaussian Mixture based Deep Generative Model) is a generative model derived from a dynamic Gaussian mixture distribution. It makes predictions based on incomplete inputs. DGM-O is instantiated with multilayer perceptron (MLP).
mTAND (Shukla and Marlin, 2021): mTAND (Multi-Time Attention network) is a deep learning model that learns representations of continuous time values and uses an attention mechanism to generate a consistent representation of a time series based on a varying amount of observations.
SeFT (Horn et al., 2020): SeFT (Set Functions for Time Series) addresses irregularly-sampled time series. It is based on differentiable set function learning.
UTDE (Zhang et al., 2023b): UTDE (Unified TDE module) is built upon TDE (Temporal discretization-based embedding). It models asynchronous time series data by combining imputation embeddings and learned interpolation embeddings through a gating mechanism. It also uses a time attention mechanism.
@@
| CNN LeCun et al. (1998) | 0.757 | 0.447 | 0.449 |
| RNN Elman (1990) | 0.750 | 0.449 | 0.433 |
| LSTM Hochreiter (1997) | 0.757 | 0.443 | 0.430 |
| Transformer Vaswani et al. (2017) | 0.749 | 0.409 | 0.433 |
| DGM-O Wu et al. (2021) | 0.671 | 0.325 | 0.382 |
| mTAND Shukla and Marlin (2019) | 0.743 | 0.437 | 0.441 |
| UTDE Zhang et al. (2023b) | 0.758 | 0.453 | 0.445 |
@@
| Flat Deznabi et al. (2021) | 0.755 | 0.447 | 0.437 |
| HierTrans Pappagari et al. (2019) | 0.754 | 0.425 | 0.434 |
| mTAND Shukla and Marlin (2019) | 0.757 | 0.435 | 0.440 |
@@
| MMTM Joze et al. (2020) | 0.769 | 0.469 | 0.459 |
| DAFT Pölsterl et al. (2021) | 0.765 | 0.452 | 0.458 |
| MedFuse Hayat et al. (2022) | 0.763 | 0.461 | 0.454 |
| DrFuse Yao et al. (2024) | 0.748 | 0.443 | 0.442 |
@@ B.2 Baselines only using clinical notes
Flat Deznabi et al. (2021): Flat encodes clinical notes using a fine-tuned BERT model. It also utilizes an LSTM model that takes in patients’ vital signals so as to jointly model the two modalities. Furthermore, it addresses the temporal irregularity issue of modeling patients’ vital signals.
HierTrans Pappagari et al. (2019): HierTrans (Hierarchical Transformers) is built upon BERT model. It achieves an enhanced ability to take in long inputs by first partitioning the inputs into shorter sequences and processing them separately. Then, it propagates each output via a recurrent layer.
T-LSTM Baytas et al. (2017): T-LSTM (Time-Aware LSTM) deals with irregular time intervals in EHRs by learning decomposed cell memory which models elapsed time. The final patient subtyping model uses T-LSTM in an auto-encoder module before doing patient subtyping.
FT-LSTM Zhang et al. (2020a): FT-LSTM (Flexible Time-aware LSTM Transformer) models the multi-level structure in clinical notes. At the base level, it uses a pre-trained ClinicalBERT model. Then, it merges sequential information and content embedding into a new position-enhanced representation. Then, it uses a time-aware layer that considers the irregularity of time intervals.
GRU-D (Che et al., 2018): Please refer to the above subsection.
mTAND (Shukla and Marlin, 2021): Please refer to the above subsection.
@@ B.3 Baselines using multimodal EHR
MMTM (Joze et al., 2020): MMTM (Multi-modal Transfer Module) is a neural network module that leverages knowledge from various modalities in CNN. It can recalibrate features in each CNN stream via excitation and squeeze operations.
DAFT (Pölsterl et al., 2021): DAFT (Dynamic Affine Feature Map Transform) is a general-purpose CNN module that alters the feature maps of a convolutional layer with respect to a patient’s clinical data.
MedFuse (Hayat et al., 2022): MedFuse is an LSTM-based fusion module capable of processing both uni-modal and multi-modal input. It treats multi-modal representations of data as a sequence of uni-modal representations. It handles inputs of various lengths via the recurrent inductive bias of LSTM.
DrFuse (Yao et al., 2024): DrFuse is a fusion module that addresses the issue of missing modality by separating the unique features within each modality and the common ones across modalities. It also adds a disease-wise attention layer for each modality.
@@
| CNN (LeCun et al., 1998) | 74.16 | 36.32 | 27.34 | 67.50 | 37.00 | 20.90 |
| RNN (Elman, 1990) | 77.40 | 38.87 | 23.93 | 66.89 | 36.28 | 18.72 |
| LSTM Hochreiter (1997) | 79.40 | 41.95 | 34.17 | 67.59 | 37.21 | 20.87 |
| Transformer (Vaswani et al., 2017) | 75.73 | 39.76 | 21.45 | 67.44 | 37.33 | 20.44 |
| IP-Net (Shukla and Marlin, 2019) | 77.58 | 46.02 | 36.8 | 68.20 | 38.4 | 20.62 |
| GRU-D (Che et al., 2018) | 53.72 | 17.54 | 8.99 | 50.53 | 23.14 | 9.42 |
| DGM-O (Wu et al., 2021) | 70.50 | 33.79 | 14.14 | 59.74 | 29.95 | 4.33 |
| mTAND (Shukla and Marlin, 2021) | 80.63 | 45.17 | 33.01 | 66.87 | 36.38 | 19.07 |
| SeFT (Horn et al., 2020) | 61.97 | 25.13 | - | 57.10 | 26.98 | - |
| UTDE (Zhang et al., 2023b) | 80.89 | 45.08 | 37.53 | 67.68 | 37.58 | 17.63 |
@@
| Flat Deznabi et al. (2021) | 62.37 | 23.97 | 23.38 | 66.10 | 36.04 | 40.40 |
| HierTrans Pappagari et al. (2019) | 61.51 | 20.96 | 17.95 | 58.36 | 27.84 | 33.61 |
| T-LSTM Baytas et al. (2017) | 53.62 | 16.82 | 21.40 | 58.56 | 28.12 | 33.06 |
| FT-LSTM Zhang et al. (2020a) | 48.63 | 15.35 | 11.85 | 54.46 | 25.29 | 29.12 |
| GRU-D (Che et al., 2018) | 56.02 | 18.63 | 23.16 | 57.94 | 27.85 | 27.13 |
| mTAND (Shukla and Marlin, 2021) | 62.80 | 24.64 | 24.82 | 68.31 | 38.53 | 40.76 |
@@
| MMTM (Joze et al., 2020) | 80.65 | 48.96 | 47.40 | 70.28 | 39.61 | 43.32 |
| DAFT (Pölsterl et al., 2021) | 81.87 | 47.79 | 48.91 | 70.87 | 40.22 | 44.10 |
| MedFuse (Hayat et al., 2022) | 81.60 | 48.35 | 48.12 | 70.82 | 40.39 | 44.03 |
| DrFuse (Yao et al., 2024) | 80.94 | 45.64 | 48.58 | 70.27 | 39.90 | 43.43 |

### v3

@@ 4.1 Experimental Setup
Evaluation Metrics. The 48-hour In-Hospital Mortality (48-IHM) prediction is a binary classification with a marked label imbalance, indicated by a death-to-discharge ratio of approximately 1:6. Following previous work (harutyunyan2019multitask; zhang2023improving), we use AUROC, AUPR, and F1 score, for a comprehensive evaluation. The -hour Phenotype Classification (24-PHE) involves predicting the presence of 25 different medical conditions during an ICU stay, making it a multi-label classification task. For this task, we employ the AUROC, AUPR, and F1 score (Macro) for a thorough assessment of model efficacy. The F1 score threshold is determined by selecting the value that maximizes the F1 score on the validation set.
Implementation Details. We train the model with batch size of 128, learning rate of 4e-5, and Adam (kingma2014adam) optimizer. We use a cosine annealing learning rate scheduler with a 0.2 warm-up proportion. To prevent overfitting, we implement early stopping when there is no increase in the AUROC on the validation set for 48-IHM or 24-PHENO over consecutive epochs. All experiments are conducted on 1 RTX-3090 GPU card using about 1 hour per run. We clip the norm of gradient values with 0.5 for stable training. By default, we use Bert-tiny DBLP:journals/corr/abs-1908-08962; bhargava2021generalization as our text encoder.
Compared Methods. To ensure a comprehensive comparison, we compare our CTPD with three types baselines: MITS-only approaches, note-only approaches and multimodal approaches. For MITS-only setting, we compare CTPD with 4 baselines for imputed regular time series: RNN (elman1990finding), LSTM (hochreiter1997long), CNN (lecun1998gradient) and Transformer vaswani2017attention, and 5 baselines for irregular time series, including IP-Net (shukla2019interpolation), GRU-D (che2018recurrent), DGM-O (wu2021dynamic), mTAND (shukla2021multi), SeFT (horn2020set), and UTDE (zhang2023improving). The imputation approach follows the MIMIC-III benchmark (harutyunyan2019multitask). For the note-only setting, we compare our model with 6 baselines: Flat deznabi2021predicting, HierTrans pappagari2019hierarchical, T-LSTM baytas2017patient, FT-LSTM zhang2020time, GRU-D (che2018recurrent), and mTAND (shukla2021multi). In the multimodal setting, we compare our model with 4 baselines: MMTM joze2020mmtm, DAFT polsterl2021combining, MedFuse hayat2022medfuse, and DrFuse yao2024drfuse. To ensure a fair comparison, we implement Bert-tiny bhargava2021generalization; DBLP:journals/corr/abs-1908-08962 as the text encoder across all baselines. Details of baselines can be found in the Appendix B.
@@ 4.2 Comparison with SOTA Baselines
Discussion on Missing Modalities and Noisy Data Scenarios. Currently, CTPD framework is built on paired time-series and clinical notes. In practice, the dataset might have missing modalities, such as partial clinical notes or time-series data are missed. We acknowledge this is an interesting research direction. However, the study of missing modalities falls outside the scope of our work and will be explored in future research. Additionally, our evaluation is conducted on MIMIC-III johnson2016mimic, a large-scale, de-identified real-world dataset containing patient records from critical care units at Beth Israel Deaconess Medical Center between 2001 and 2012. Our approach is designed to be adaptable and can be seamlessly applied to other real-world databases.
@@ 6 Limitations
A limitation of our approach is its primary focus on extracting temporal semantics in the embedding space, which affects model interpretability. As a potential solution, we plan to explore retrieval-based methods patel2024retrieve or pretrained generative models zhao2023rleg to enhance the interpretability of learned temporal pattern embeddings in future work. Besides, our work primarily focuses on exploring cross-modal temporal patterns to enhance representation learning, the scenario of missing modalities falls outside the current scope—since shared patterns across modalities cannot be fully captured in their absence. We plan to thoroughly investigate this direction in future work. Additionally, our framework is currently designed for specific clinical prediction tasks. In practice, there are various prediction tasks related to multimodal EHR analysis. Extending the proposed method into a more generalized or foundational model capable of handling multiple downstream tasks with minimal training annotations could be more practical and effective. Our future work will focus on resolving those aspects.
@@ Appendix A Experimental Setup Details
We presented the class distributions for 48-IHM and 24-PHE in Table 7 and Table 8, respectively. 48-IHM is a binary classification task, whereas 24-PHE is a multi-label classification problem with 25 labels. Notably, our 24-PHE task differs from the phenotype classification problem in the MIMIC-III benchmark but follows the setup in zhang2023improving. This setup focuses on acute care conditions that arise during ICU stays, where early prediction is critical for timely intervention. To enhance clinical relevance, we use only the first 24 hours of data for phenotype classification rather than the entire admission record. As highlighted in yang2021multimodal, early-stage diagnosis holds greater clinical significance.
The selection of 25 phenotype labels for the 24-PHE task follows established practices in the MIMIC-III benchmark harutyunyan2019multitask, as also utilized in prior studies like UTDE zhang2023improving and MedFuse hayat2022medfuse. These labels cover conditions commonly observed in adult ICUs, including 12 critical and life-threatening conditions (e.g., respiratory failure, sepsis), 8 chronic conditions that are often considered comorbidities or risk factors (e.g., diabetes, metabolic disorders), and 5 ‘mixed’ conditions that exhibit characteristics of both chronic and acute conditions. Phenotype labels were determined using the MIMIC-III ICD-9 diagnosis table. To facilitate the translation and conversion of the abovementioned conditions, we use the Health Cost and Utilization (HCUP) Clinical Classification Software (CCS)222https://hcup-us.ahrq.gov/toolssoftware/ccs/ccs.jsp. We first mapped each ICD-9 code to its corresponding HCUP CCS category, retaining only 25 categories. Diagnoses were then linked to ICU stays using the hospital admission identifier, as ICD-9 codes in MIMIC-III are associated with hospital visits rather than specific ICU stays. To reduce label ambiguity, we excluded hospital admissions involving multiple ICU stays, ensuring each diagnosis could be associated with a single ICU stay. Please find the class distribution of 25 phenotypes in Table 8.
@@ B.1 Baselines only using MITS
CNN (lecun1998gradient): CNN (Convolutional Neural Network) uses backpropagation to synthesize a complex decision surface that facilitates learning.
RNN (elman1990finding): RNN (Residual Neural Network) is trained to process data sequentially so as to model the time dimension of data.
LSTM hochreiter1997long: LSTM (Long short-term memory) is a variant of recurrent neural network. It excels at dealing with the vanishing gradient problem and is relatively insensitive to time gap length.
Transformer (vaswani2017attention): Transformer is a powerful deep learning architecture based on attention mechanism. It has great generalisability and has been adopted as foundation model in multiple research domains.
IP-Net (shukla2019interpolation): IP-Net (Interpolation-Prediction Network) is a deep learning architecture for supervised learning focusing on processing sparse multivariate time series data that are sampled irregularly.
GRU-D (che2018recurrent): GRU-D is based on GRU (Gated Recurrent Unit). It incorporates features of missing data in EHR into the model architecture to improve prediction results.
DGM-O (wu2021dynamic): DGM-O (Dynamic Gaussian Mixture based Deep Generative Model) is a generative model derived from a dynamic Gaussian mixture distribution. It makes predictions based on incomplete inputs. DGM-O is instantiated with multilayer perceptron (MLP).
mTAND (shukla2021multi): mTAND (Multi-Time Attention network) is a deep learning model that learns representations of continuous time values and uses an attention mechanism to generate a consistent representation of a time series based on a varying amount of observations.
SeFT (horn2020set): SeFT (Set Functions for Time Series) addresses irregularly-sampled time series. It is based on differentiable set function learning.
UTDE (zhang2023improving): UTDE (Unified TDE module) is built upon TDE (Temporal discretization-based embedding). It models asynchronous time series data by combining imputation embeddings and learned interpolation embeddings through a gating mechanism. It also uses a time attention mechanism.
@@
| CNN lecun1998gradient | 0.757 | 0.447 | 0.449 |
| RNN elman1990finding | 0.750 | 0.449 | 0.433 |
| LSTM hochreiter1997long | 0.757 | 0.443 | 0.430 |
| Transformer vaswani2017attention | 0.749 | 0.409 | 0.433 |
| DGM-O wu2021dynamic | 0.671 | 0.325 | 0.382 |
| mTAND shukla2019interpolation | 0.743 | 0.437 | 0.441 |
| UTDE zhang2023improving | 0.758 | 0.453 | 0.445 |
@@
| Flat deznabi2021predicting | 0.755 | 0.447 | 0.437 |
| HierTrans pappagari2019hierarchical | 0.754 | 0.425 | 0.434 |
| mTAND shukla2019interpolation | 0.757 | 0.435 | 0.440 |
@@
| MMTM joze2020mmtm | 0.769 | 0.469 | 0.459 |
| DAFT polsterl2021combining | 0.765 | 0.452 | 0.458 |
| MedFuse hayat2022medfuse | 0.763 | 0.461 | 0.454 |
| DrFuse yao2024drfuse | 0.748 | 0.443 | 0.442 |
@@ B.2 Baselines only using clinical notes
Flat deznabi2021predicting: Flat encodes clinical notes using a fine-tuned BERT model. It also utilizes an LSTM model that takes in patients’ vital signals so as to jointly model the two modalities. Furthermore, it addresses the temporal irregularity issue of modeling patients’ vital signals.
HierTrans pappagari2019hierarchical: HierTrans (Hierarchical Transformers) is built upon BERT model. It achieves an enhanced ability to take in long inputs by first partitioning the inputs into shorter sequences and processing them separately. Then, it propagates each output via a recurrent layer.
T-LSTM baytas2017patient: T-LSTM (Time-Aware LSTM) deals with irregular time intervals in EHRs by learning decomposed cell memory which models elapsed time. The final patient subtyping model uses T-LSTM in an auto-encoder module before doing patient subtyping.
FT-LSTM zhang2020time: FT-LSTM (Flexible Time-aware LSTM Transformer) models the multi-level structure in clinical notes. At the base level, it uses a pre-trained ClinicalBERT model. Then, it merges sequential information and content embedding into a new position-enhanced representation. Then, it uses a time-aware layer that considers the irregularity of time intervals.
GRU-D (che2018recurrent): Please refer to the above subsection.
mTAND (shukla2021multi): Please refer to the above subsection.
@@ B.3 Baselines using multimodal EHR
MMTM (joze2020mmtm): MMTM (Multi-modal Transfer Module) is a neural network module that leverages knowledge from various modalities in CNN. It can recalibrate features in each CNN stream via excitation and squeeze operations.
DAFT (polsterl2021combining): DAFT (Dynamic Affine Feature Map Transform) is a general-purpose CNN module that alters the feature maps of a convolutional layer with respect to a patient’s clinical data.
MedFuse (hayat2022medfuse): MedFuse is an LSTM-based fusion module capable of processing both uni-modal and multi-modal input. It treats multi-modal representations of data as a sequence of uni-modal representations. It handles inputs of various lengths via the recurrent inductive bias of LSTM.
DrFuse (yao2024drfuse): DrFuse is a fusion module that addresses the issue of missing modality by separating the unique features within each modality and the common ones across modalities. It also adds a disease-wise attention layer for each modality.
@@
| CNN (lecun1998gradient) | 74.16 | 36.32 | 27.34 | 67.50 | 37.00 | 20.90 |
| RNN (elman1990finding) | 77.40 | 38.87 | 23.93 | 66.89 | 36.28 | 18.72 |
| LSTM hochreiter1997long | 79.40 | 41.95 | 34.17 | 67.59 | 37.21 | 20.87 |
| Transformer (vaswani2017attention) | 75.73 | 39.76 | 21.45 | 67.44 | 37.33 | 20.44 |
| IP-Net (shukla2019interpolation) | 77.58 | 46.02 | 36.8 | 68.20 | 38.4 | 20.62 |
| GRU-D (che2018recurrent) | 53.72 | 17.54 | 8.99 | 50.53 | 23.14 | 9.42 |
| DGM-O (wu2021dynamic) | 70.50 | 33.79 | 14.14 | 59.74 | 29.95 | 4.33 |
| mTAND (shukla2021multi) | 80.63 | 45.17 | 33.01 | 66.87 | 36.38 | 19.07 |
| SeFT (horn2020set) | 61.97 | 25.13 | - | 57.10 | 26.98 | - |
| UTDE (zhang2023improving) | 80.89 | 45.08 | 37.53 | 67.68 | 37.58 | 17.63 |
@@
| Flat deznabi2021predicting | 62.37 | 23.97 | 23.38 | 66.10 | 36.04 | 40.40 |
| HierTrans pappagari2019hierarchical | 61.51 | 20.96 | 17.95 | 58.36 | 27.84 | 33.61 |
| T-LSTM baytas2017patient | 53.62 | 16.82 | 21.40 | 58.56 | 28.12 | 33.06 |
| FT-LSTM zhang2020time | 48.63 | 15.35 | 11.85 | 54.46 | 25.29 | 29.12 |
| GRU-D (che2018recurrent) | 56.02 | 18.63 | 23.16 | 57.94 | 27.85 | 27.13 |
| mTAND (shukla2021multi) | 62.80 | 24.64 | 24.82 | 68.31 | 38.53 | 40.76 |
@@
| MMTM (joze2020mmtm) | 80.65 | 48.96 | 47.40 | 70.28 | 39.61 | 43.32 |
| DAFT (polsterl2021combining) | 81.87 | 47.79 | 48.91 | 70.87 | 40.22 | 44.10 |
| MedFuse (hayat2022medfuse) | 81.60 | 48.35 | 48.12 | 70.82 | 40.39 | 44.03 |
| DrFuse (yao2024drfuse) | 80.94 | 45.64 | 48.58 | 70.27 | 39.90 | 43.43 |